OpenStudy (anonymous):

4-5/8 X 4 how does this come to 3/2 and break this down step by step

OpenStudy (hoblos):

4 - 5/8 x 4 = 4 - 5/2 = 8/2 - 5/2 =3/2

OpenStudy (anonymous):

So i subtract first?

OpenStudy (hoblos):

no you multiply and sorry for the previous one :D

OpenStudy (anonymous):

I did this for you on your other post. I will break it down even further if that is what you need. You have to solve the multiplication first. \[4 - [5/8*4]\] \[[5/8*4/1]\] you get 20/8 4 = 32/8 you need to find the least common multiple in order to add or subtract fractions so, you end up with (32/8) - (20/8) = 12/8 This can be reduced down to 3/2 by a factor of 4. Hope this helps.

OpenStudy (anonymous):

can someone help e with mine?

OpenStudy (anonymous):

THANKS SO MUCH

OpenStudy (anonymous):

No problem, please click good answer to give me a medal. Thanks!